OpenStudy (anonymous):

The gas piston of an M4 carbine rifle with in initial position of x0=5[mm].experiences a velocity defined by x’=3+5t for one second. What is the piston’s finale position?

OpenStudy (anonymous):

plz solve it in details ...thanks

OpenStudy (experimentx):

\[ {dx \over dt} = 3 + 5t \\ \int dx = \int( 3 + 5t) dt = x = 5t^2/2 + 3t + c \\ put, x=0, when \;t = 0 \\ that \; would \; give \; c =5, \\ now \; put \; t=1 \] in 5t^2/2 + 3t + 5 and get your answer.

OpenStudy (anonymous):

wat is t^2 ???

OpenStudy (experimentx):

\[ {5t^2 \over 2} + 3t + 5 \]

OpenStudy (anonymous):

yea tanx aloooooooooooooooooooooot it's 10.5 mm